. The length of a given cylindrical wire is increased by 100%  Due to the consequent decrease in diameter the change in the resistance of the wire will be
 
(a) 200%
(b) 100%
(c) 50%
(d) 300%

Grade:6

3 Answers

bharat bajaj IIT Delhi
9 years ago
As we know
R = p L/A
New l` = 2l ..
Due to increase in length, area must decrease as volume remains constant..
Al = A`l`
Hence, A` = A/2
New Resistance R` = pl`/A` = 4pl/A
Change in resistance = 3pl/A or 300 %
D) is the answer
THanks
Bharat Bajaj
IIT Delhi
askiitians faculty
Navjyot Kalra
9 years ago

(d) The total volume remains the same before and after stretching.

Therefore A x l = A’ x l’

Here l’ = 2 l

A’ = A x l / l’ = A x l / 2 l = A/2

Percentage change in resistance

= Rf – Ri / Rl x 100 = ? l’/A’ - l/A / l ? l/A x 100

= [ (l’/A’ x A/ l) – 1 ] x 100 [(2 l/A/2 x A / l)-1 ] x 100

= 300%

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

As we know
R = p L/A
New l` = 2l ..
Due to increase in length, area must decrease as volume remains constant..
Al = A`l`
Hence, A` = A/2
New Resistance R` = pl`/A` = 4pl/A
Change in resistance = 3pl/A or 300 %

Thanks and Regards